Halliford Corporation expects to have earnings this coming year of

$ 2.773$2.773

per share. Halliford plans to retain all of its earnings for the next two years.​ Then, for the subsequent two​ years, the firm will retain

54 %54%

of its earnings. It will retain

21 %21%

of its earnings from that point onward. Each​ year, retained earnings will be invested in new projects with an expected return of

21.3 %21.3%

per year. Any earnings that are not retained will be paid out as dividends. Assume​ Halliford's share count remains constant and all earnings growth comes from the investment of retained earnings. If​ Halliford's equity cost of capital is

8.3 %8.3%​,

what price would you estimate for Halliford​ stock?
Answer #1

Halliford’s dividend forecast (g = retention rate × return on new investment)

Year

0

1

2

3

4

5

6

Earnings

1).EPS Growth Rate (vs. prior yr)

21.30%

21.30%

11.502%

11.502%

4.473%

2). EPS

$2.773

$3.3636

$4.080

$4.5494

$5.0727

$5.2996

Dividends

3). Retention Ratio

100%

100%

54%

54%

21%

21%

4). Dividend Payout Ratio

0%

0%

46%

46%

79%

79%

Dividends(2 x 4) $0 $0    $1.8768 $2.0927 $4.0074 $4.1867

From year 5 on, dividends grow at constant rate of 4.473%. Therefore,

P(5) = D6 / (r - gC) = $4.1867 / (8.3% - 4.473%) = $4.1867 / 0.03827 = $109.40

P(0) = [D3 / (1 + r)3] + [D4 / (1 + r)4] + [(D5 + P5) / (1 + r)5]

= [$1.8768 / 1.0833] + [$2.0927 / 1.0834] + [($4.0074 + $109.40) / 1.0835]

= $1.48 + $1.52 + $76.12 = $79.12